# Enabling Local Neural Operators to perform Equation-Free System-Level Analysis

### v1

@@
[…]
Enabling Local Neural Operators
to perform Equation-Free System-Level Analysis
Gianluca Fabiani Hopkins Extreme Materials Institute, Johns Hopkins University, Baltimore, 21218, MD, USA Dept. of Chemical and Biomolecular Engineering, Johns Hopkins University, Baltimore, 21218, MD, USA Hannes Vandecasteele Dept. of Chemical and Biomolecular Engineering, Johns Hopkins University, Baltimore, 21218, MD, USA Dept. of Applied Mathematics and Statistics, Johns Hopkins University, Baltimore, 21218, MD, USA Somdatta Goswami Dept. of Civil and Systems Engineering, Johns Hopkins University, Baltimore, 21218, MD, USA Constantinos Siettos Dipartimento di Matematica e Applicazioni “Renato Caccioppoli”, Università degli Studi di NapoliFederico II, Naples, Italy Corresponding authors, emails: constantinos.siettos@unina.it, yannisk@jhu.edu Ioannis G. Kevrekidis Dept. of Chemical and Biomolecular Engineering, Johns Hopkins University, Baltimore, 21218, MD, USA Dept. of Applied Mathematics and Statistics, Johns Hopkins University, Baltimore, 21218, MD, USA School of Medicine’s Dept. of Urology, Johns Hopkins University, Baltimore, 21218, MD, USA Corresponding authors, emails: constantinos.siettos@unina.it, yannisk@jhu.edu
{onecolabstract}
@@
Computer-assited modeling of dynamical systems is fundamental to understanding, predicting, and analyzing the behavior of real-world processes across science and engineering. Based on such models, and their system-level analysis – which extends beyond “brute-force” temporal simulations – we can enable physics-based decision-making, management, and control of the dynamics [1, 2, 3, 4]. Key objectives include computing long-term steady states / limit cycles (even unstable ones that are unreachable through direct temporal simulations) and constructing bifurcation diagrams to detect and understand the mechanisms that trigger qualitative dynamic transitions, including catastrophic (“hard”, tipping point) bifurcations [2, 1, 5, 6, 7]. Such computer-aided system-level analysis is crucial to tackle real-world challenges, including runaway reactions, pandemic outbreaks, climate tipping points, economic crises, ecosystem extinctions, cancer progression, and neurological disorders [1, 8, 9, 10, 11, 12].
@@
Neural Operators (NOs) have been introduced to address the latter problem by learning mappings between function spaces. Various architectures for NOs have been proposed over the past few years: Deep Operator Networks (DeepONets) [23, 24, 25], Fourier Neural Operators (FNOs) [26], Graph Neural Operators (GNOs) [27, 28], Wavelet Neural Operators (WNOs) [29], Spectral Neural Operators (SNOs) [30] and more recently Random Projection-based Operator Networks (RandONets) [31]. However, while NOs learn mappings for infinite-dimensional PDEs, they still face the challenges related to the “curse of dimensionality”. In fact, NOs built using DNNs typically suffer from slow convergence because their optimization is NP-hard [32], and their overparameterization makes training extremely computationally expensive [33, 34, 35, 36, 37, 38]. Hence, despite their theoretical promise [39, 40, 41], DNNs and NOs based on them often provide only partially satisfactory numerical approximation accuracy, which in practice hinder their efficiency for tasks requiring high-precision numerical analysis. Recently, to tackle some of the computational challenges mentioned above, we introduced RandONet [31], a randomized neural operator architecture. RandONets share the same fundamental structure as DeepONets – both rooted in the Chen and Chen framework [41] – but replace trainable hidden layers in DNNs with just one hidden layer (as in [41]), using fixed randomized embeddings, drawing on random projections, leveraging the Johnson-Lindenstrauss lemma [42] and the universal approximation power of random features [43, 44, 45, 46, 47, 7, 48, 31]. This design transforms the original non-convex, gradient-based (and computational expensive) training used in DNNs into a linear system for the output weights. Although the resulting system may be high-dimensional and ill-conditioned, it can be solved effectively using standard numerical methods such as the Moore–Penrose pseudoinverse, Tikhonov regularization [49], QR decomposition, or complete orthogonal decomposition (COD) [50].
@@
Here, for the first time, we integrate NOs with the EF framework, enabling system-level analysis based on data-driven approximations of PDE solution operators. This extends the role of NOs beyond simple simulations to construct bifurcation diagrams and detect tipping points in high-dimensional systems. The framework is, in principle, applicable to any NO architecture; its practical effectiveness will, of course, depend on the accuracy and generalization performance of the trained local neural operator it is built around. Our key idea is to train local NOs - defined over small time steps, and possibly in small space-time patches – and then use them as building blocks of computational workflows inspired by classical numerical analysis schemes. To illustrate the proposed framework, we use three representative examples: (a) the Allen-Cahn phase-field PDE with homogeneous Neumann boundary conditions [51], a system that undergoes consecutive pitchfork bifurcations; (b) the Liouville-Bratu-Gelfand PDE [52] which exhibits a saddle-node bifurcation; and (c) the FitzHugh-Nagumo model [5, 53, 7], a pair of coupled reaction-diffusion PDEs that have Hopf and saddle-node bifurcations. We show that the proposed approach results in high-accuracy predictions in identifying critical points and constructing solution branches, even unstable ones, robustly confirming the correct stability.
@@
Here, we initially focus on learning a data-driven approximation of the so-called “time-stepper”/solution operator of PDEs using NOs. The learned time-stepper provides a global solver-free surrogate of the PDE, mapping input functions (e.g., initial conditions and parameters) to the system’s state at a future time. Then the learned time-stepper, can be used iteratively for system-level computations (see e.g. [54], and more below about Projective Integration). In the standard multiscale EF framework, this timestepper is computed on demand, every time its application is required, and then discarded; in our EF Local NO framework, the same trained neural timestepper is used with different initial/boundary condition inputs as necessary.
@@
Regarding the availability of data for the output function $v$, there are two scenarios. In the first scenario, the value ($v(\bm{y}_{i})$) of functions in the output are known at a fixed grid ${\bm{y}_{1},\bm{y}_{2},\dots,\bm{y}_{n}}$, where $y_{i}\in Y$. This case is termed as ‘aligned’ data. In the second scenario, the output grid may be sparse and limited, vary randomly, or come from multiple experiments with different grid sizes, leading to what is known as ‘unaligned’ data [55]. In both cases, the “trunk” network is responsible for processing the spatial locations of the output data. This is a case of particular interest, and we will focus on this introducing “Gap-Tooth” NOs, a novel approach tailored to this context.
@@
Performing accurate bifurcation analysis with NOs, including the reliable detection of critical points, is a particularly challenging task. A simple approach that is currently in use is the construction of bifurcation diagrams via long temporal simulations of the NOs. Toward this aim, as noted in [54], it is often more efficient to learn the system’s evolution, $\mathcal{S}_{\Delta t}$ over shorter time steps $\Delta t$, and then to autoregressively apply the learned time stepper via
@@
To mitigate this issue, an effective approach is to employ a fixed-point iteration method, such as Newton-Raphson, to directly find the steady-state solutions as fixed points of the short time solution operator, bypassing the need for long-time simulations. This approach mitigates error accumulation in steady-state predictions and, by constructing appropriate initial conditions, enables the convergence to unstable steady states that are unreachable via simple temporal simulations. A schematic overview of this Equation-Free strategy, including the Neural Operator time-stepper, finite-difference Jacobian approximation, and matrix-free Newton-Krylov and Arnoldi iterative methods, is shown in Fig. 1.
@@
A full bifurcation diagram as a function of the parameter $\lambda$ can then be computed using pseudo-arclength continuation [22]. In matrix-free numerical analysis, the stability of steady states can be assessed using the Arnoldi algorithm [19, 18], which approximates the dominant eigenpairs relevant to stability. We refer to Appendix section E for more details. At this point, we note that, as highlighted in [57], the discrete-time solution operator $\mathcal{S}_{\Delta t}$ has favorable numerical properties: its discrete nature inherently leads to a more compact and controlled spectrum, compared to the continuous counterpart, where the eigenvalues are often confined to a region that is easier to analyze and manage. In particular, the spectrum of the Jacobian matrix of the time-stepper tends to be more tightly grouped, leading to better conditioning and more stable numerical behavior. This property allows for improved convergence rates in linear iterations, such as those used in pseudo-arclength continuation, and can facilitate the derivation of more accurate bounds on convergence. This makes stability and bifurcation analysis in discrete-time operators more robust [57].
@@
Although our primary focus is on performing accurate bifurcation and stability analysis, multiscale computations can enhance the applicability of NOs for accelerating computations and learning more efficiently from sparse and limited data. Two notable techniques in this context are (Telescopic) Projective Integration (PI) and the ‘Gap-Tooth‘ Scheme, initially developed in the multiscale EF framework [3, 58, 59, 60], that we briefly discuss here. Both exploit the smoothness of PDE solutions in time (PI) and in space (Gaptooth).
@@
Given the current state $u_{n}$, at the $n$-th projective integration step, we first apply the time-stepper (the short time NO) for $k$ small steps
@@
which are separated by $N-1$ inactive intervals (gaps) of size typically $\Delta\xi\geq\Delta x$. At each time step $t=0,\Delta t,2\Delta t,\dots,T$, we interpolate the boundary values, $u\big{(}x_{i}-\frac{\Delta x}{2},s\big{)}$ and $u\big{(}x_{i}+\frac{\Delta x}{2},s\big{)}$, for each tooth $T_{i}$, using computed solutions in neighboring teeth $T_{i-1}$ and $T_{i+1}$, for $s\in[t,t+\Delta t]$, estimating the values of the right and left flux (boundary) $F^{right}_{i}$ and $F^{left}_{i}$. Then we learn (and compute) the local-tooth ($T_{i}$) solution operator $\mathcal{S}_{\Delta t}^{T_{i}}[u_{s},F^{right}_{i},F^{left}_{i}]$ within each tooth $T_{i}$ over $s\in[t,t+\Delta t]$ using these interpolated boundary conditions ($F^{right}_{i},F^{left}_{i}$), as additional branch-inputs for the Gap-tooth NOs. When there is no explicit spatial dependence, the learned local NOs are identical (same weights and biases) for each tooth, that is, $\mathcal{S}_{\Delta t}^{T_{i}}\equiv\mathcal{S}_{\Delta t}^{T_{j}}$, $\forall i,j=1,\dots,N$. This enables a “patch” learning process that reduces computational complexity.
Both PI and Gap-Tooth methods offer promising ways to enhance the efficiency of NO-based solvers, particularly for multiscale systems where fine-scale resolution is computationally expensive. Most importantly, the two methods can be combined – we can have local-in-space and local-in-time NOs; these trained NOs will be used over only part of the spatial domain (taking advantage of the solution smoothness in space-time). This is the so-called “patch dynamics” EF scheme, which now becomes a “patch NO” scheme for large space-time, even geometry agnostic domains.
@@ Results
We present numerical results for three benchmark problems focused on identifying the solution operator of parameter-dependent PDEs: (a) the Allen-Cahn PDE; (b) a parabolic extension of the Liouville–Bratu–Gelfand PDE; (c) a Lattice-Boltzmann implementation of the FitzHugh-Nagumo (FHN) PDE. In particular, as described in the methodology, we assessed the performance of our local NOs to reconstruct the corresponding bifurcation diagrams, using Newton-GMRES and pseudo-arclength continuation to converge on steady states (stable or unstable) and track unstable solution branches past turning points. We also attempted to approximate the correct dominant eignespectrum using a local NO-based Arnoldi method. Furthermore, we explored the use of the ‘Gap-Tooth” NOs: local in space solution operators, aiming for greater data efficiency and for increased adaptability to changing domains and boundary conditions. For our first (the Allen-Cahn) example, we perform a direct comparison between two training strategies for NOs: the standard gradient-based approach for the fully-trained DeepONet variant, and the random projection-based RandONets variant. While both models share architectural structure features, their different training algorithms lead to distinct generalization behavior. In our own experiments, the “standard” DeepONet was sometimes only partially successful in reconstructing the full bifurcation diagram, occasionally missing solution branches, or producing spurious ones. The RandONet formulation was more consistently accurate, recovering both solution branches as well as their dominant eigenspectrum. For the other two case studies we present the results obtained just with RandONets (DeepONet computations were also performed, consistent with the above observations). The important issue here, in our opinion, is not to focus on contrasting the relative performance of various NO implementations in various examples; it is rather to demonstrate practically that all implementations can benefit from our Local NO framework.
@@
(e) partial and spurious bifurcaiton diagram
Figure 2: DeepONet-based timestepper for the Allen-Cahn PDE in Eq. (14), trained with data with multiple parameters in the range $\varepsilon\in[0.22,0.7]$. The results are compared with the approximated solutions, obtained using Finite-Difference (FD) discretization in space and forward Euler in time of the known PDE. (a) Time evolution for a sine-type initial condition for $\varepsilon=0.22$; (b) corresponding pointwise absolute error; (c) sine-type steady-state solution for $\varepsilon=0.22$; (d) inaccurate eigenvalue spectrum of Jacobian computed for the steady-state; (e) arc-length continuation of DeepONet steady-states give rise to an incomplete and spurious bifurcation diagram.
@@
with $\Omega=[-1,1]$ and $0<\varepsilon\ll 1$ is a parameter controlling the interface width. The phase field variable $\phi(x,t)$ represents two distinct phases (a value of $+1$ representing one phase, and $-1$ representing the other). The function $W(\phi)=\frac{1}{4}\left(\phi^{2}-1\right)^{2}$ is a double-well potential forcing $\phi$ to assume values close to ±$\pm 1$ to clearly differentiate the phases and to make the boundaries between them narrow. The Allen-Cahn PDE arises naturally as the $L^{2}$-gradient flow of the Ginzburg-Landau free energy functional,
[…]
@@
$-\varepsilon\phi_{xx}+\frac{1}{\varepsilon}\left(\phi^{3}-\phi\right)=0,\quad-1<x<1,$ (16)
with Neumann BCs: $\phi_{x}(-1)=\phi_{x}(1)=0$.
@@
Initial conditions, to generate data, are sampled from a family of functions that capture the different types of transient dynamics, eventually converging to the different multiple steady-state solutions. These include sine, cosine, and flat states, randomized by cosine-modulated Gaussian profiles. The datasets are further structured to ensure symmetry. Full details on data generation and preprocessing are provided in Appendix section B.1.
@@
In our DeepONet implementation, the neural operator consists of $6$ hidden layers of $100$ neurons in both the branch and trunk nets, followed by an additional fully connected layer with $300$ neurons, yielding 7 hidden layers in total for both trunk and branch nets. Thus, the training process consists in optimizing more than $200,000$ trainable parameters. The implementation is carried out in Python using the JAX backend, ensuring compatibility with automatic differentiation and GPU acceleration. Training is performed using the Adam optimizer with a learning rate of $10^{-3}$, for 100,000 epochs. The training process takes approximately $3.91\times 10^{3}$ seconds (about 1 hour) on a high-performance GPU cluster (ARCH core facility at Johns Hopkins University and the Rockfish cluster). The resulting model achieves a mean squared error (MSE) of $1.60\times 10^{-7}$ on the test dataset. Finally, after the training of DeepONet and RandONets, we attempted to reconstruct the full bifurcation diagram, using Newton-GMRES and pseudo-arclength continuation. Both the DeepONet and RandONet time steppers are compared against a reference solver based on finite differences in space and a forward Euler method in time. For numerical stability, the Euler solver uses a fixed time step of $dt=0.00005$. In contrast, DeepONet and RandONet were trained to approximate dynamics with a larger time step of $dt=0.01$, which is $200$ times larger. As a result, despite the initial training cost, DeepONet and RandONet enable significantly faster simulations. Finally, to enforce the symmetry required in pitchfork bifurcation, we used an explicit transformation of the DeepONet/RandONet operator
@@
To assess the performance of the RandONet timesteppers, we first report the results obtained when RandONets were trained at a fixed value of the parameter (for our illustrations for $\varepsilon=0.22$). Then, we report the results obtained when the homotopy-based RandONets were trained to learn the dynamics in the parameter range $\varepsilon\in(0,22,0.7)$ and the corresponding bifurcation diagram obtained. Figure 3 shows the results obtained at the fixed value of the bifurcation parameter. The top row focuses on the sine-type solutions, whereas the bottom row shows the cosine-type solutions. Figures 3a and 3c display the steady-state solutions of the respective methods obtained through Newton-GMRES iterations. In Figures 3b and 3d, we present the full eigenvalue spectrum of the steady-state Jacobian of the actual PDE time-stepper. Finally, the zoom-in of Figures 3b and 3d show the three leading eigenvalues calculated by the Arnoldi method. All these figures highlight a close agreement between the RandONet and the reference solver. These results demonstrate the ability of RandONets to effectively capture both steady-state solutions and critical spectral features of the Allen-Cahn equation. Finally, to construct the bifurcation diagram, including the first two of the many pitchfork bifurcations given by equations (17) and (18), we integrated the RandONet timestepper within the pseudo-arclength continuation algorithm coupled with Newton-GMRES. The two right-most pitchfork bifurcations computed for both timesteppers are shown in Figure 4a. As can be seen, there is excellent visual agreement between the bifurcation diagrams of the actual PDE and the one constructed via the RandONet. We also computed the absolute values of the three leading eigenvalues with respect to the bifurcation parameter. The sine-type solutions are shown in Figure 4b, the cosine-type in Figure 4c, and the zero-type in Figure 4d. These results demonstrate that RandONets also accurately approximate the actual key eigenvalues, i.e., those governing the dominant dynamics and bifurcation behavior, thus accurately estimating the correct stability of the associated solution branches.

### v2

@@
Abstract
@@
Computer-assisted modeling of dynamical systems is fundamental to understanding, predicting, and analyzing the behavior of real-world processes across science and engineering. Based on such models, and their system-level analysis – which extends beyond “brute-force” temporal simulations – we can enable physics-based decision-making, management, and control of the dynamics [1, 2, 3, 4]. Key objectives include computing long-term steady states / limit cycles (even unstable ones that are unreachable through direct temporal simulations) and constructing bifurcation diagrams to detect and understand the mechanisms that trigger qualitative dynamic transitions, including catastrophic (“hard”, tipping point) bifurcations [2, 1, 5, 6, 7]. Such computer-aided system-level analysis is crucial to tackle real-world challenges, including runaway reactions, pandemic outbreaks, climate tipping points, economic crises, ecosystem extinctions, cancer progression, and neurological disorders [1, 8, 9, 10, 11, 12].
@@
Neural Operators (NOs) have been introduced to address the latter problem by learning mappings between function spaces. Various architectures for NOs have been proposed over the past few years: Deep Operator Networks (DeepONets) [23, 24, 25], Fourier Neural Operators (FNOs) [26], Graph Neural Operators (GNOs) [27, 28], Wavelet Neural Operators (WNOs) [29], Spectral Neural Operators (SNOs) [30] and more recently Random Projection-based Operator Networks (RandONets) [31]. However, while NOs learn mappings for infinite-dimensional PDEs, they still face the challenges related to the “curse of dimensionality”. In fact, NOs built using DNNs typically suffer from slow convergence because their optimization is NP-hard [32], and their overparameterization makes training extremely computationally expensive [33, 34, 35, 36, 37, 38]. Hence, despite their theoretical promise [39, 40, 41], DNNs and NOs based on them often provide only partially satisfactory numerical approximation accuracy, which in practice hinder their efficiency for tasks requiring high-precision numerical analysis. Recently, to tackle some of the computational challenges mentioned above, we introduced RandONet [31], a randomized neural operator architecture. RandONets share the same fundamental structure as DeepONets – both rooted in the Chen and Chen framework [41] – but replace trainable hidden layers in DNNs with just one hidden layer (as in [41]), using fixed randomized embeddings, drawing on random projections, leveraging the Johnson-Lindenstrauss lemma [42] and the universal approximation power of random features [43, 44, 45, 46, 47, 7, 48, 31]. This design transforms the original non-convex, gradient-based (and computationally expensive) training used in DNNs into a linear system for the output weights. Although the resulting system may be high-dimensional and ill-conditioned, it can be solved effectively using standard numerical methods such as the Moore–Penrose pseudoinverse, Tikhonov regularization [49], QR decomposition, or complete orthogonal decomposition (COD) [50].
@@
Here, for the first time, we integrate NOs with the EF framework, enabling system-level analysis based on data-driven approximations of PDE solution operators. This extends the role of NOs beyond simple simulations to construct bifurcation diagrams and detect tipping points in high-dimensional systems. The framework is, in principle, applicable to any NO architecture; its practical effectiveness will, of course, depend on the accuracy and generalization performance of the trained local neural operator it is built around. Our key idea is to train local NOs - defined over small time intervals, and possibly in small space-time patches - and then use them as modular building blocks of computational workflows inspired by classical numerical analysis schemes. Our notion of locality is distinct from recent work on Neural Operators with localized kernels [51, 52], where ‘local’ refers to the compactly supported kernels/filters used within the neural operator. In contrast, we employ local operators restricted in time or space–time, and compose them modularly into larger workflows. Related approaches include short-time Neural Operators [53, 54], which are applied iteratively to reconstruct long-time predictions. Similarly, we apply local operators iteratively, but also introduce Projective Integration (PI) [55] (and possibly, in the future, Telescopic-PI [56]) to further accelerate long-time simulations. This modularity echoes existing work on patch dynamics and space-time multiscale computation [3, 57, 58, 59], as well as recent developments on learning homogenized and effective dynamics from microscopic models [60, 61]. Related advances include the enhancement of neural operators through finite element methods and Schwarz-type domain decomposition techniques [62, 63, 64]. We envision that combining such ideas with local training can significantly broaden the range of NOs applicable to coarse-grained, equation-free modeling in complex multiscale systems. To illustrate the proposed framework, we use three representative examples: (a) the Allen-Cahn phase-field PDE with homogeneous Neumann boundary conditions [65], a system that undergoes consecutive pitchfork bifurcations; (b) the Liouville-Bratu-Gelfand PDE [66] which exhibits a saddle-node bifurcation; and (c) the FitzHugh-Nagumo model [5, 67, 7], a pair of coupled reaction-diffusion PDEs that have Hopf and saddle-node bifurcations. We show that the proposed approach results in high-accuracy predictions in identifying critical points and constructing solution branches, even unstable ones, robustly confirming the correct stability.
@@
Here, we initially focus on learning a data-driven approximation of the so-called “time-stepper”/solution operator of PDEs using NOs. The learned time-stepper provides a global solver-free surrogate of the PDE, mapping input functions (e.g., initial conditions and parameters) to the system’s state at a future time. Then the learned time-stepper, can be used iteratively for system-level computations (see e.g. [54, 53], and more below about Projective Integration). In the standard multiscale EF framework, this timestepper is computed on demand, every time its application is required, and then discarded; in our EF Local NO framework, the same trained neural timestepper is used with different initial/boundary condition inputs as necessary.
@@
Regarding the availability of data for the output function $v$, there are two scenarios. In the first scenario, the value ($v(\bm{y}_{i})$) of functions in the output are known at a fixed grid $\bm{y}_{1},\bm{y}_{2},\dots,\bm{y}_{n}$}, where $y_{i}\in Y$. This case is termed as ‘aligned’ data. In the second scenario, the output grid may be sparse and limited, vary randomly, or come from multiple experiments with different grid sizes, leading to what is known as ‘unaligned’ data [68]. In both cases, the “trunk” network is responsible for processing the spatial locations of the output data. This is a case of particular interest, and we will focus on this by introducing “Gap-Tooth” NOs, a novel approach tailored to this context.
@@
Performing accurate bifurcation analysis with NOs, including the reliable detection of critical points, is a particularly challenging task. A simple approach that is currently in use is the construction of bifurcation diagrams via long temporal simulations of the NOs. Toward this aim, as noted in [54, 53], it is often more efficient to learn the system’s evolution, $\mathcal{S}_{\Delta t}$ over shorter time steps $\Delta t$, and then to autoregressively apply the learned time stepper via
@@
To mitigate this issue, an effective approach is to employ a fixed-point iteration method, such as Newton-Raphson, to directly find the steady-state solutions as fixed points of the short time solution operator, bypassing the need for long-time simulations. This approach mitigates error accumulation in steady-state predictions and, by constructing appropriate initial conditions, enables the convergence to unstable steady states that are unreachable via simple temporal simulations. A schematic overview of this Equation-Free strategy, including the Neural Operator time-stepper, finite-difference Jacobian approximation, and matrix-free Newton-Krylov and Arnoldi iterative methods, is shown in Figure 1.
@@
A full bifurcation diagram as a function of the parameter $\lambda$ can then be computed using pseudo-arclength continuation [22]. In matrix-free numerical analysis, the stability of steady states can be assessed using the Arnoldi algorithm [19, 18], which approximates the dominant eigenpairs relevant to stability. We refer to Appendix section E for more details. At this point, we note that, as highlighted in [70], the discrete-time solution operator $\mathcal{S}_{\Delta t}$ has favorable numerical properties: its discrete nature inherently leads to a more compact and controlled spectrum, compared to the continuous counterpart, where the eigenvalues are often confined to a region that is easier to analyze and manage. In particular, the spectrum of the Jacobian matrix of the time-stepper tends to be more tightly grouped, leading to better conditioning and more stable numerical behavior.
Figure 2: Schematic representation of (a) The Euler timestepper; (b) Projective Integration applied for NOs; (c) The Gap-Tooth scheme applied to an NO; and (d) The full patches scheme for NOs. Regions shaded in gray indicate where the timestepper or neural operator is active, white regions is where they are not. Dashed arrows indicate communication of boundary conditions through local polynomial interpolation in the Gap-Tooth and Patches schemes.
This property allows for improved convergence rates in linear iterations, such as those used in pseudo-arclength continuation, and can facilitate the derivation of more accurate bounds on convergence. This makes stability and bifurcation analysis in discrete-time operators more robust [70].
@@
Although our primary focus is on performing accurate bifurcation and stability analysis, multiscale computations can enhance the applicability of NOs for accelerating computations and learning more efficiently from sparse and limited data.
Two notable techniques in this context are (Telescopic) Projective Integration (PI) and the ‘Gap-Tooth‘ Scheme, initially developed in the multiscale EF framework [3, 55, 56, 71], that we briefly discuss here. Both exploit the smoothness of PDE solutions in time (PI) and in space (Gap-Tooth). The combination of PI and Gap-Tooth leads to the so-called patch dynamics [57], for which we have also proposed patch Neural Operators (patch NOs). A schematic representation of PI, Gap-Tooth, and their combination in patch dynamics is shown in Figure 2.
@@
Given the current state $u_{n}$, at the $n$-th projective integration step, we first apply the time-stepper (the short time NO) for $k$ small steps,
@@
which are separated by N1N-1 inactive intervals (gaps) of size typically $\Delta\xi\geq\Delta x$. At each time step $t=0,\Delta t,2\Delta t,\dots,T$, we interpolate the boundary values, $u\big(x_{i}-\frac{\Delta x}{2},s\big)$ and $u\big(x_{i}+\frac{\Delta x}{2},s\big)$, for each tooth $T_{i}$, using computed solutions in neighboring teeth $T_{i-1}$ and $T_{i+1}$, for $s\in[t,t+\Delta t]$, estimating the values of the right and left flux (boundary) $F^{right}_{i}$ and $F^{left}_{i}$. Then we learn (and compute) the local-tooth ($T_{i}$) solution operator $\mathcal{S}_{\Delta t}^{T_{i}}[u_{s},F^{right}_{i},F^{left}_{i}]$ within each tooth $T_{i}$ over $s\in[t,t+\Delta t]$ using these interpolated boundary conditions ($F^{right}_{i},F^{left}_{i}$), as additional branch-inputs for the Gap-tooth NOs. When there is no explicit spatial dependence, the learned local NOs are identical (same weights and biases) for each tooth, that is, $\mathcal{S}_{\Delta t}^{T_{i}}\equiv\mathcal{S}_{\Delta t}^{T_{j}}$, $\forall i,j=1,\dots,N$. This enables a “patch” learning process that reduces computational complexity.
Both PI and Gap-Tooth methods offer promising ways to enhance the efficiency of NO-based solvers, particularly for multiscale systems where fine-scale resolution is computationally expensive. Most importantly, the two methods can be combined – we can have local-in-space and local-in-time NOs; these trained NOs will be used over only part of the spatial domain (taking advantage of the solution smoothness in space-time). This is the so-called “patch dynamics” EF scheme [57], which now becomes a “patch NO” scheme for large space-time, even geometry agnostic domains.
@@ Results
We present numerical results for three benchmark problems focused on identifying the solution operator of parameter-dependent PDEs: (a) the Allen-Cahn PDE; (b) a parabolic extension of the Liouville–Bratu–Gelfand PDE; (c) a Lattice-Boltzmann implementation of the FitzHugh-Nagumo (FHN) PDE. In particular, as described in the methodology, we assessed the performance of our local NOs to reconstruct the corresponding bifurcation diagrams, using Newton-GMRES and pseudo-arclength continuation to converge on steady states (stable or unstable) and track unstable solution branches past turning points. We also attempted to approximate the correct dominant eigenspectrum using a local NO-based Arnoldi method. Furthermore, we explored the use of the “Gap-Tooth” NOs: local in space solution operators, aiming for greater data efficiency and for increased adaptability to changing domains and boundary conditions. For our first (the Allen-Cahn) example, we perform a direct comparison between two training strategies for NOs: the standard gradient-based approach for the fully-trained DeepONet variant, and the random projection-based RandONets variant. While both models share architectural structure features, their different training algorithms lead to distinct generalization behavior. In our own experiments, the “standard” DeepONet was sometimes only partially successful in reconstructing the full bifurcation diagram, occasionally missing solution branches, or producing spurious ones. The RandONet formulation was more consistently accurate, recovering both solution branches as well as their dominant eigenspectrum. For the other two case studies we present the results obtained just with RandONets (DeepONet computations were also performed, consistent with the above observations). The important issue here, in our opinion, is not to focus on contrasting the relative performance of various NO implementations in various examples; it is rather to demonstrate practically that all implementations can benefit from our Local NO framework.
@@
(e) partial and spurious bifurcation diagram
Figure 3: DeepONet-based timestepper for the Allen-Cahn PDE in Eq. (14), trained with data with multiple parameters in the range $\varepsilon\in[0.22,0.7]$. The results are compared with the approximated solutions, obtained using Finite-Difference (FD) discretization in space and forward Euler in time of the known PDE. (a) Time evolution for a sine-type initial condition for $\varepsilon=0.22$; (b) corresponding point-wise absolute error; (c) sine-type steady-state solution for $\varepsilon=0.22$; (d) inaccurate eigenvalue spectrum of Jacobian computed for the steady-state; (e) arc-length continuation of DeepONet steady-states give rise to an incomplete and spurious bifurcation diagram.
@@
with $\Omega=[-1,1]$ and $0<\varepsilon\ll 1$ is a parameter controlling the interface width. The phase field variable $\phi(x,t)$ represents two distinct phases (a value of +1+1 representing one phase, and 1-1 representing the other). The function $W(\phi)=\frac{1}{4}\left(\phi^{2}-1\right)^{2}$ is a double-well potential forcing $\phi$ to assume values close to $\pm 1$ to clearly differentiate the phases and to make the boundaries between them narrow. The Allen-Cahn PDE arises naturally as the $L^{2}$-gradient flow of the Ginzburg-Landau free energy functional,
$E(\phi)=\int_{-1}^{1}\left(\frac{\varepsilon}{2}\phi_{x}^{2}+\frac{1}{4\varepsilon}\left(\phi^{2}-1\right)^{2}\right)dx,$ (15)
@@
$\varepsilon\phi_{xx}+\frac{1}{\varepsilon}\left(\phi^{3}-\phi\right)=0,\quad-1<x<1,$ (16)
with Neumann boundary conditions: $\phi_{x}(-1)=\phi_{x}(1)=0$.
@@
Initial conditions, to generate data, are sampled from a family of functions that capture the different types of transient dynamics, eventually converging to the different multiple steady-state solutions. These include sine, cosine, and flat states, randomized by cosine-modulated Gaussian profiles. The datasets are further structured to ensure symmetry. For training and short-time predictions, we used a local time step of $\Delta t=0.01$. For longer horizon predictions, the neural operator was applied autoregressively. Full details on data generation and preprocessing are provided in Appendix section B.1.
@@
In our DeepONet implementation, the neural operator consists of $6$ hidden layers of $100$ neurons in both the branch and trunk nets, followed by an additional fully connected layer with $300$ neurons, yielding 7 hidden layers in total for both trunk and branch nets. Thus, the training process consists of optimizing more than $200,000$ trainable parameters. The implementation is carried out in Python using the JAX backend, ensuring compatibility with automatic differentiation and GPU acceleration. Training is performed using the Adam optimizer with a learning rate of $10^{-3}$, for 100,000 epochs. The training process takes approximately $3.91\times 10^{3}$ seconds (about 1 hour) on a high-performance GPU cluster (ARCH core facility at Johns Hopkins University and the Rockfish cluster). The resulting model achieves a mean squared error (MSE) of $1.60\times 10^{-7}$ on the test dataset. Finally, after the training of DeepONet and RandONets, we attempted to reconstruct the full bifurcation diagram, using Newton-GMRES and pseudo-arclength continuation. Both the DeepONet and RandONet time steppers are compared against a reference solver based on finite differences in space and a forward Euler method in time. For numerical stability, the Euler solver uses a fixed time step of $dt=0.00005$. In contrast, DeepONet and RandONet were trained to approximate dynamics with a larger time step of $dt=0.01$, which is $200$ times larger. As a result, despite the initial training cost, DeepONet and RandONet enable significantly faster simulations. Finally, to enforce the symmetry required in pitchfork bifurcation, we used an explicit transformation of the DeepONet/RandONet operator
@@
To assess the performance of the RandONet timesteppers, we first report the results obtained when RandONets were trained at a fixed value of the parameter (for our illustrations for $\varepsilon=0.22$). Then, we report the results obtained when the homotopy-based RandONets were trained to learn the dynamics in the parameter range $\varepsilon\in(0,22,0.7)$ and the corresponding bifurcation diagram obtained. Figure 4 shows the results obtained at the fixed value of the bifurcation parameter. The top row focuses on the sine-type solutions, whereas the bottom row shows the cosine-type solutions. Figures 4a and 4c display the steady-state solutions of the respective methods obtained through Newton-GMRES iterations. In Figures 4b and 4d, we present the full eigenvalue spectrum of the steady-state Jacobian of the actual PDE time-stepper. Finally, the zoom-in of Figures 4b and 4d shows the three leading eigenvalues calculated by the Arnoldi method. All these figures highlight a close agreement between the RandONet and the reference solver. These results demonstrate the ability of RandONets to effectively capture both steady-state solutions and critical spectral features of the Allen-Cahn equation. Finally, to construct the bifurcation diagram, including the first two of the many pitchfork bifurcations given by equations (17) and (18), we integrated the RandONet timestepper within the pseudo-arclength continuation algorithm coupled with Newton-GMRES. The two right-most pitchfork bifurcations computed for both timesteppers are shown in Figure 5a. As can be seen, there is excellent visual agreement between the bifurcation diagrams of the actual PDE and the one constructed via the RandONet. We also computed the absolute values of the three leading eigenvalues with respect to the bifurcation parameter. The sine-type solutions are shown in Figure 5b, the cosine-type in Figure 5c, and the zero-type in Figure 5d. These results demonstrate that RandONets also accurately approximate the actual key eigenvalues, i.e., those governing the dominant dynamics and bifurcation behavior, thus accurately estimating the correct stability of the associated solution branches.